OpenStudy (anonymous):

∫∫ x sqrt(y^2-x^2) dA, R={(x,y)|0≤x≤y, 1≤x≤2}

OpenStudy (luigi0210):

This: \[\LARGE \int\limits(\int\limits~x~\sqrt{y^2-x^2})dA\]

OpenStudy (luigi0210):

Wait, are those the limits on the right?

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\int\limits_{0}^{y} x \sqrt{y ^{2}+x ^{2}}dx dy\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (shamil98):

luigi do you know fubini's theorem?

OpenStudy (anonymous):

the answer is \[5(2\sqrt{2}-1)/4\]

OpenStudy (awkwardpanda):

Knew it.

OpenStudy (anonymous):

please show me the solution

OpenStudy (shamil98):

oh I thought luigi asked this... mb

OpenStudy (awkwardpanda):

wait no.. I'm sorry I was kidding.. ;-;

OpenStudy (anonymous):

lol

OpenStudy (shamil98):

R={(x,y)|0≤x≤y, 1≤x≤2} Integrate with respect to y first.

OpenStudy (usukidoll):

what type is it? type one or type two? type one is dydx with this drawing of a vertical line and the limits are on the horizontal x axis type two is dxdy with a drawing of a horizontal line and the limits are on the veritical y axis

OpenStudy (usukidoll):

well we can tell according to the limits...if it's all numbers it's usually at the outer integral so in this case we have a dxdy so type 2

OpenStudy (usukidoll):

because if we integrate with respect to y first we have a problem...you shouldn't have any variables

OpenStudy (usukidoll):

if we switch the problem around we need the new limits of integration... sadly my calculus iv professor umm sucks hardcore...

OpenStudy (usukidoll):